OpenStudy (spring98):

I need help asap. Will medal and fan! Thanks! The mass of a freight train is 9.98 • 10^6 kilograms. The mass of an aircraft carrier is 7.3 • 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (spring98):

@dan815

OpenStudy (spring98):

no one seems to know these questions. Can you help?

OpenStudy (spring98):

@Shalante

OpenStudy (spring98):

@shadow13

OpenStudy (spring98):

@luckycoins888 @west_coast @Elsa213 @razor99 @taylor1048 @Meehan98 @dan815 plz help me I need help.

OpenStudy (spring98):

this is scientific notation

OpenStudy (spring98):

@countryboy01

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

6.62 times 10^5

OpenStudy (anonymous):

ok im getting 25 i might be wrong but im not sure

OpenStudy (spring98):

this is supposed to be in scientific notation

OpenStudy (spring98):

@dan815

OpenStudy (anonymous):

im not sure then i got it messed up

OpenStudy (anonymous):

im sorry

OpenStudy (jdoe0001):

\(\large \begin{array}{lrlll} 7.3\times 10^7\implies &7\underline{3000000}\\ 9.98\times 10^6\implies &-9\underline{980000} \\\hline\\ &? \end{array}\)

OpenStudy (spring98):

@jdoe0001 is "Approximately, 63.02 *10^6 more kilograms is the mass of an aircraft carrier than the mass of a freight train" the answer?

OpenStudy (jdoe0001):

yes that's correct BUT scientific notation, uses only one number to the left of the dot you have two numbers \(\bf \underline{63}.02\times 10^6\)

OpenStudy (spring98):

so is it 6.3?

OpenStudy (jdoe0001):

\(\bf 6.302\times 10^7\) yeap

OpenStudy (jdoe0001):

same thing, we simply move the dot over to the left one more slot so we end up with only one digit on the left side :)

OpenStudy (spring98):

thanks

OpenStudy (jdoe0001):

yw